Fes Company is making adjusting journal entries for the year ended December 31, 2018. In developing information for the adjusting journal entries, you learned the following: A two-year insurance premium of $6,900 was paid on January 1, 2018, for coverage beginning on that date. As of December 31, 2018, the unadjusted balances were $6,900 for Prepaid Insurance and $0 for Insurance Expense. At December 31, 2018, you obtained the following data relating to supplies. Unadjusted balance in Supplies on December 31, 2018 $ 13,500 Unadjusted balance in Supplies Expense on December 31, 2018 69,000 Supplies on hand, counted on December 31, 2018 8,800

Required:

Of the $6,900 paid for insurance, what amount should be reported on the 2018 income statement as Insurance Expense?

What amount should be reported on the December 31, 2018, balance sheet as Prepaid Insurance?

What amount should be reported on the 2018 income statement as Supplies Expense?

What amount should be reported on the December 31, 2018, balance sheet as Supplies? Indicate the accounting equation effects of the adjustment required for (a) insurance and (b) supplies.
Answer #1

Insurance expense = 6900/2 = 3450 Prepaid expense = 6900-3450 = 3450 2) Supplies consumed or expenses = (unadjusted supplies)Assets Liabilities + Stockholders equity Prepaid insurance Insurance expense -3450 -3450 B Supplies Supplies expense -4400 -